# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ Abstract
We develop novel LASSO-based methods for coefficient testing and confidence interval construction in the Gaussian linear model with $n\geq d$. Our methods’ finite-sample guarantees are identical to those of their ubiquitous ordinary-least-squares-$t$-test-based analogues, yet have substantially higher power when the true coefficient vector is sparse. In particular, our coefficient test, which we call the $\ell$-test, performs like the one-sided $t$-test (despite not being given any information about the sign) under sparsity, and the corresponding confidence intervals are more than 10% shorter than the standard $t$-test based intervals. The nature of the $\ell$-test directly provides a novel exact adjustment conditional on LASSO selection for post-selection inference, allowing for the construction of post-selection p-values and confidence intervals. None of our methods require resampling or Monte Carlo estimation. We perform a variety of simulations and a real data analysis on an HIV drug resistance data set to demonstrate the benefits of the $\ell$-test. We end with a discussion of how the $\ell$-test may asymptotically apply to a much more general class of parametric models.
@@ 1.1 Motivation
where $\bm{X}\in\mathbb{R}^{n\times d}$ is full column-rank (and in particular, assume $n\geq d$) and treated as non-random, and $\bm{\beta}\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}_{>0}$ are unknown. For a given covariate of interest $X_{j}$ this paper will consider testing $H_{j}:\beta_{j}=0$ and the related problem of constructing a confidence interval for $\beta_{j}$. It will leverage the LASSO to do so and our method’s construction will also make it easy to construct conditionally valid versions, conditioned on LASSO selection.
The go-to solution for this type of single covariate inference is based on the linear regression $t$-test for $H_{j}$, which can be efficiently inverted to obtain a $t$-test-based confidence interval for $\beta_{j}$. The linear regression $t$-test dates back over a century (Fisher, 1922) and is ubiquitous in introductory statistics courses and methods courses in nearly every domain of science and engineering. As a result, it is hard to overstate how universally widely used it is in practice. And it is easy to see why: the $t$-test is intuitive, easy to compute, and comes with strong theoretical guarantees.
The goal of this paper is to allow an analyst to leverage a belief in sparsity (of $\bm{\beta}$) to conduct more informative inference (when sparsity holds) without sacrificing the statistical guarantees of the $t$-test (even when sparsity does not hold). Sparsity is a widely held belief throughout applications in science and engineering (indeed, this belief is so ubiquitous that it has a name: the principle of parsimony), and while leveraging sparsity in regression is a heavily studied subject (we review existing approaches in Section 1.3), methods developed to leverage it often rely on sparsity for both validity and increased power, while we explicitly seek to rely on it only for increased power.
@@ 1.2 Summary of our contributions
We develop a hypothesis test for $H_{j}:\beta_{j}=0$, which we call the $\ell$-test, that uses the absolute value of the fitted LASSO coefficient as its test statistic. Using novel analysis of the conditional distribution of the LASSO estimator given the sufficient statistic of the linear model, we derive the test statistic’s exact null distribution, allowing us to efficiently compute p-values without resampling or Monte Carlo. We argue that the $\ell$-test will often achieve nearly the power of the one-sided $t$-test when $\bm{\beta}$ is sparse, despite not knowing the sign of $\beta_{j}$ and remaining exactly valid under identical assumptions as the two-sided $t$-test. We show the $\ell$-test can be efficiently inverted to produce exact confidence intervals, which due to the power improvement of the $\ell$-test are typically more than about 10% shorter than $t$-test-based confidence intervals when $\bm{\beta}$ is sparse. For both the $\ell$-test and its corresponding confidence interval, we show that a cross-validation procedure can be used to select the penalty parameter $\lambda$ in the LASSO from the data without impacting our validity guarantees, making our proposed procedures tuning-parameter-free. The nature of the $\ell$-test and our formula for its null distribution make it straightforward to derive and compute (novel) post-selection $\ell$-test p-values such that the post-selection $\ell$-test p-value for $H_{j}$ is exactly (and non-conservatively) valid conditional on the LASSO estimate of $\beta_{j}$ being nonzero; this conditional test can also be inverted to construct a confidence interval that is valid conditional on selection. A wide range of simulations and an application to HIV drug resistance demonstrate our methods to be powerful, efficient, and robust. In our discussion, we point out that the $\ell$-test can be directly generalized to any normal means problem with known covariance, and hence in particular should be applicable to any parametric model’s maximum likelihood estimator in its asymptotic Gaussian limit.
@@ 1.3 Related work
The standard choice for testing $H_{j}:\beta_{j}=0$ (and, via inversion, constructing confidence intervals) in the linear model is the (two-sided) $t$-test, or the one-sided $t$-test when the sign is known. Given the age and ubiquity of the linear model in statistics, we do not attempt to cover all related literature (though Lei and Bickel (2020), particularly their Appendix B, provides an excellent and detailed review), but just note that many such works focus on developing methods with some form of guarantees under weaker assumptions than the standard (homoskedastic) Gaussian linear model assumed in this paper (Friedman, 1937, Pitman, 1937, 1938, Kruskal and Wallis, 1952, Tukey, 1958, Hajek, 1962, Adichie, 1967, Jaeckel, 1972, Efron, 1979, Freedman, 1981, Gutenbrunner et al., 1993, Lei and Bickel, 2020). As their goal is robustness, these methods generally do not outperform the $t$-test in the (homoskedastic) Gaussian linear model, whereas this is exactly the goal of the current paper: maintain the same guarantees as the $t$-test while improving its power when $\bm{\beta}$ is sparse. To our knowledge, the only other work that leverages a linear model’s sparsity for testing an individual coefficient is the de-biased LASSO (van de Geer et al., 2014, Zhang and Zhang, 2014, Javanmard and Montanari, 2014), but, when $n\geq d$, it either reduces exactly to the $t$-test or is only asymptotically valid under strong sparsity assumptions on $\bm{\beta}$; in contrast, the validity of the $\ell$-test holds regardless of the sparsity of $\bm{\beta}$. There are a number of excellent works that aim to leverage sparsity for more powerful inference in the Gaussian linear model, including (fixed-X) knockoffs (Barber and Candès, 2015) and subsequent follow-ups that improve its performance (e.g., Spector and Janson (2022), Luo et al. (2022), Ren and Barber (2023), Lee and Ren (2024)), as well as methods based on mirror statistics (Xing et al., 2023, Dai et al., 2022), but all of these methods can only be used for variable selection and do not provide single-variable inference.
@@
Two approaches share a similar goal as ours in seeking to improve the $t$-test’s power without making further assumptions. The first approach is Habiger and Peña (2014), which proposes to split the observations into two disjoint (thus independent) parts and uses the first part to estimate the sign of $\beta_{j}$ and then leverages the estimated sign in a test using the second part of the data. As we will discuss in Section 2.2, the $\ell$-test also leverages an estimated sign of $\beta_{j}$, but since it does not involve data splitting, it does not suffer from the associated loss in sample size and power. The second approach is called Frequentist, assisted by Bayes (FAB), which, given a prior on $\bm{\beta}$ and $\sigma^{2}$, produces Bayes-optimal power (or confidence interval width) subject to maintaining the same frequentist validity as the $t$-test (Hoff and Yu, 2019). But this paper only considers (dense) Gaussian priors for $\bm{\beta}$ and hence, unlike the $\ell$-test, does not leverage sparsity. [^1] The discussion section of Hoff and Yu (2019) mentions the possibility of using FAB with spike-and-slab priors to incorporate sparsity but does not pursue it.
@@
While the goal of the $\ell$-test is most similar to the works mentioned so far, its approach is most closely related to the idea of conditioning on a sufficient statistic under the null hypothesis. This approach is perhaps most prominently used in constructing uniformly most powerful unbiased tests (Lehmann and Scheffé, 1955), including the $t$-test. But this idea is also fundamental to co-sufficient sampling (Bartlett, 1937, Stephens, 2012), which is used for testing in a wide variety of contexts; see Barber and Janson (2022) for a recent review of such tests. However, the $\ell$-test is not sampling-based, and besides, to the best of our knowledge, the only work applying co-sufficient sampling to testing in the linear model is Huang and Janson (2020), but there it is applied to knockoffs (Barber and Candès, 2015, Candès et al., 2018) which is a method for variable selection and cannot perform single-coefficient inference.
@@
Related to this paper’s post-selection inference methods, there is a rich literature on obtaining p-values that are valid post-selection (Cox, 1975, Berk et al., 2013, Fithian et al., 2014, Rinaldo et al., 2016, Bachoc et al., 2016, Tian and Taylor, 2018) and in particular in the linear model conditioned on LASSO selection (Lee et al., 2016, Tibshirani et al., 2016, Liu et al., 2018, Panigrahi et al., 2021). Closest to our work is Liu et al. (2018), which can be thought of as adjusting the standard $t$-test (or, more accurately, the $z$-test, since they assume $\sigma^{2}$ known) for a coefficient to make it valid conditional on that coefficient being selected by the LASSO. The conditional $\ell$-test can be thought of as making an analogous adjustment to the $\ell$-test rather than the $t$-test, resulting in similar gains under sparsity over Liu et al. (2018)’s method as the unconditional $\ell$-test achieves over the standard $t$-test; see Sections 4, 5.4, and Appendix F.4 for further comparison and discussion.
@@ 1.5 Software
Functions in R (R Core Team, 2024) for running all methods in this paper are available at github.com/SSouhardya/l-test.
@@ 2.1 The $\ell$-distribution
As the first step towards characterizing the $\ell$-distribution $F^{\lambda}_{\ell}(\cdot\mid\bm{S}^{(j)})$, we restate (Luo et al., 2022, Proposition E.1) that exactly characterizes the distribution of $\bm{y}\mid\bm{S}^{(j)}$ under $H_{j}$. Let $\bm{P}_{-j}=\bm{X}_{-j}(\bm{X}_{-j}^{T}\bm{X}_{-j})^{-1}\bm{X}_{-j}^{T}$ denote the projection matrix onto the column space of $\bm{X}_{-j}$.
Lemma 2.1 (Luo et al. (2022)).
@@
where $\mathbb{S}^{n-d}$ denotes the unit sphere of dimension $n-d$.
For completeness, a proof of the lemma is provided in Section C.1 of the Appendix. Next our main theoretical result, Theorem 2.1, establishes a mapping between $\hat{\beta}_{j}^{\lambda}$ and just the first element of $\bm{u}$, $u_{1}$ (there is nothing special about index 1 here except that we defined $\bm{V}$ to have only its first column non-orthogonal to $\bm{X}_{j}$), which will give an immediate characterization of $F_{\ell}^{\lambda}$ via the known distribution of $u_{1}$ from Equation (2.2).
Theorem 2.1 (Characterization of the $\ell$-distribution).
Consider the unique decomposition (2.1) from Lemma 2.1 and for any $b\in\mathbb{R}$ and $\epsilon\in\{-1,1\}$, define the functions
@@
Theorem 2.1 exactly characterizes the $\ell$-distribution: for $b\neq 0$, we have
@@
Thus, in particular, if we let $F_{u}$ denote the CDF of $u_{1}$ under $H_{j}$, then $F_{\ell}^{\lambda}(b)=\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}\leq b\mid\bm{S}^{(j)})=F_{u}(\Lambda_{j}(b,\mathrm{sign}(b)))$ because $\Lambda_{j}(b,\mathrm{sign}(b))$ is a function of $\bm{S}^{(j)}$ for any fixed $b$. Similarly, for $b=0$, it follows from Theorem 2.1 that $F_{\ell}^{\lambda}(0)=F_{u}\left(\Lambda_{j}(0,1)\right)$. Note that $F_{u}$ is easily evaluated via a one-to-one mapping to a $t$-distribution, namely, $\frac{\sqrt{n-d}u_{1}}{\sqrt{1-u_{1}^{2}}}\sim t_{n-d}$ (see Appendix C.5 for a proof), which, along with the relations above, can be used to explicitly calculate quantiles of the $\ell$-distribution. The proof of Theorem 2.1 in Appendix C.3 hinges on two main ideas—first, we use blockwise coordinate descent to characterize the event $\{\hat{\beta}_{j}^{\lambda}=b\}$ in terms of $u_{1}$, and second, we characterize $f_{\bm{S}^{(j)}}$ by obtaining an exact expression for $\frac{\partial}{\partial u_{1}}f_{\bm{S}^{(j)}}$, which turns out to be non-negative throughout, thereby showing $f_{\bm{S}^{(j)}}$ is non-decreasing. We will see next that Theorem 2.1 also provides critical insights into the power of the $\ell$-test.

### v2

@@ Abstract
We develop novel LASSO-based methods for coefficient testing and confidence interval construction in the Gaussian linear model with $n\geq d$. Our methods’ finite-sample validity is identical to that of their ubiquitous ordinary-least-squares-$t$-test-based analogues, yet have substantially higher power when the true coefficient vector is sparse. In particular, under sparsity our coefficient test, which we call the $\ell$-test, performs like the one-sided $t$-test (despite not being given any information about the sign), and $\ell$-test-based confidence intervals are correspondingly shorter than the standard $t$-test-based intervals. The nature of the $\ell$-test directly provides a novel exact adjustment conditional on LASSO selection for post-selection inference, allowing for the construction of post-selection $p$-values and confidence intervals. None of our methods require resampling or Monte Carlo estimation. We perform a variety of simulations and a real data analysis on an HIV drug resistance data set to demonstrate the benefits of the $\ell$-test. We additionally show that the $\ell$-test can be applied to a large class of asymptotically Gaussian estimators, dramatically expanding its applicability beyond linear models.
@@ 1.1 Motivation
where $\bm{X}\in\mathbb{R}^{n\times d}$ is full column-rank (and in particular, assume $n\geq d$) and treated as non-random, and $\bm{\beta}\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}_{>0}$ are unknown. For a given covariate of interest $X_{j}$ this paper will consider testing $H_{j}:\beta_{j}=0$ and the related problem of constructing a confidence interval for $\beta_{j}$. It will leverage the LASSO (LASSO) to do so and our method’s construction will also make it easy to construct conditionally valid versions, conditioned on LASSO selection.
The go-to solution for this type of single covariate inference is based on the linear regression $t$-test for $H_{j}$, which can be efficiently inverted to obtain a $t$-test-based confidence interval for $\beta_{j}$. The linear regression $t$-test dates back over a century (fisher1922goodness) and is ubiquitous in introductory statistics courses and methods courses in nearly every domain of science and engineering. As a result, it is hard to overstate how universally widely used it is in practice. And it is easy to see why: the $t$-test is intuitive, easy to compute, and comes with strong theoretical guarantees.
The goal of this paper is to allow an analyst to leverage a belief in sparsity (of $\bm{\beta}$) to conduct more informative inference (when sparsity holds) without sacrificing the statistical guarantees of the $t$-test (even when sparsity does not hold). Sparsity is a widely held belief throughout applications in many fields, as evidenced by the widespread adoption of the LASSO (LASSO) and other sparse regression methods; see rish2014sparse, 64524c4c-35f2-309c-807b-12f074681130 for a textbook and review paper, respectively, on sparse regression in general, including discussion of applications, and, e.g., Lucas_Carvalho_Wang_Bild_Nevins_West_2006, annurev:/content/journals/10.1146/annurev-economics-061109-080451, https://doi.org/10.1002/cem.1418 for reviews on the important role of sparsity specifically in the domains of genomics (Section 6.5 will demonstrate our methods on a concrete data set in this domain), economics, and chemistry, respectively. While leveraging sparsity in regression is a heavily studied subject (we review existing approaches in Section 1.3), methods developed to leverage it often rely on sparsity for both validity and increased power, while we explicitly seek to rely on it only for increased power.
@@ 1.2 Summary of our contributions
We develop a hypothesis test for $H_{j}:\beta_{j}=0$, which we call the $\ell$-test, that uses the absolute value of the fitted LASSO coefficient as its test statistic. Using novel analysis of the conditional distribution of the LASSO estimator given the sufficient statistic of the linear model, we derive the test statistic’s exact null distribution, allowing us to efficiently compute $p$-values without resampling or Monte Carlo. We argue that the $\ell$-test will often achieve nearly the power of the one-sided $t$-test when $\bm{\beta}$ is sparse, despite not knowing the sign of $\beta_{j}$ and remaining exactly valid under identical assumptions as the two-sided $t$-test. We show the $\ell$-test can be efficiently inverted to produce exact confidence intervals, which due to the power improvement of the $\ell$-test are substantially shorter than the $t$-test-based confidence intervals when $\bm{\beta}$ is sparse; in our simulations, we find the typical improvement under sparsity to be about $10\%$. For both the $\ell$-test and its corresponding confidence interval, we show that a cross-validation procedure can be used to select the penalty parameter $\lambda$ in the LASSO from the data without impacting our validity guarantees, making our proposed procedures tuning-parameter-free. The nature of the $\ell$-test and our formula for its null distribution make it straightforward to derive and compute (novel) post-selection $\ell$-test $p$-values such that the post-selection $\ell$-test $p$-value for $H_{j}$ is exactly (and non-conservatively) valid conditional on the LASSO estimate of $\beta_{j}$ being nonzero; this conditional test can also be inverted to construct a confidence interval that is valid conditional on selection. We also show that the $\ell$-test can be extended to the rich class of locally asymptotically normal models, yielding asymptotically valid $p$-values. Furthermore, we leverage the framework of Approximate Message Passing (AMP) to perform a theoretical power analysis in medium dimensions that sheds analytical light on the power benefits of the $\ell$-test. A wide range of simulations and an application to HIV drug resistance demonstrate our methods to be powerful, efficient, and robust.
@@ 1.3 Related work
The standard choice for testing $H_{j}:\beta_{j}=0$ (and, via inversion, constructing confidence intervals) in the linear model is the (two-sided) $t$-test, or the one-sided $t$-test when the sign is known. Given the age and ubiquity of the linear model in statistics, we do not attempt to cover all related literature (though leibickel2020, particularly their Appendix B, provides an excellent and detailed review), but just note that many such works focus on developing methods with some form of guarantees under weaker assumptions than the standard (homoskedastic) Gaussian linear model assumed in this paper (friedman1937, pitman1937b, 1938, krushkalwallis1952, tukey1958, hajek1962, adichie1967, jaeckel1972, efron1979, freedman1981, gutenbrunner1993, leibickel2020). As their goal is robustness, these methods generally do not outperform the $t$-test in the (homoskedastic) Gaussian linear model, whereas this is exactly the goal of the current paper: maintain the same guarantees as the $t$-test while improving its power when $\bm{\beta}$ is sparse. To our knowledge, the only other work that leverages a linear model’s sparsity for testing an individual coefficient is the de-biased LASSO (debiasedLASSO, zhangandzhang2014, JMLR:v15:javanmard14a), but, when $n\geq d$, it either reduces exactly to the $t$-test or is only asymptotically valid under strong sparsity assumptions on $\bm{\beta}$; in contrast, the validity of the $\ell$-test holds regardless of the sparsity of $\bm{\beta}$. There are a number of excellent works that aim to leverage sparsity for more powerful inference in the Gaussian linear model, including (fixed-X) knockoffs (knockoff) and subsequent follow-ups that improve its performance (e.g., AS-LJ:2020, luo2022, renandbarber(2023), lee2024boosting), as well as methods based on mirror statistics (gaussmirror, junfdrdatasplit), but all of these methods can only be used for variable selection and do not provide single-variable inference.
@@
Two approaches share a similar goal as ours in seeking to improve the $t$-test’s power without making further assumptions. The first approach is HABIGER2014153, which proposes to split the observations into two disjoint (thus independent) parts and uses the first part to estimate the sign of $\beta_{j}$ and then leverages the estimated sign in a test using the second part of the data. As we will discuss in Section 2.2, the $\ell$-test also leverages an estimated sign of $\beta_{j}$, but since it does not involve data splitting, it does not suffer from the associated loss in sample size and power. The second approach is called Frequentist, assisted by Bayes (FAB), which, given a prior on $\bm{\beta}$ and $\sigma^{2}$, produces Bayes-optimal power (or confidence interval width) subject to maintaining the same frequentist validity as the $t$-test (yuandhofflinearregression). But this paper only considers (dense) Gaussian priors for $\bm{\beta}$ and hence, unlike the $\ell$-test, does not leverage sparsity [^2] The discussion section of yuandhofflinearregression mentions the possibility of using FAB with spike-and-slab priors to incorporate sparsity but does not pursue it.
@@
While the goal of the $\ell$-test is most similar to the works mentioned so far, its approach is most closely related to the idea of conditioning on a sufficient statistic under the null hypothesis. This approach is perhaps most prominently used in constructing uniformly most powerful unbiased tests (lehmannscheffe1955), including the $t$-test. But this idea is also fundamental to co-sufficient sampling (bartlett1937properties, CSS), which is used for testing in a wide variety of contexts; see RB-LJ:2020 for a recent review of such tests. However, the $\ell$-test is not sampling-based, and besides, to the best of our knowledge, the only work applying co-sufficient sampling to testing in the linear model is huang2020, but there it is applied to knockoffs (knockoff, EC-ea:2018) which is a method for variable selection and cannot perform single-coefficient inference.
@@
Related to this paper’s post-selection inference methods, there is a rich literature on obtaining $p$-values that are valid post-selection (Cox1975ANO, berk2013, fithianthesis, bootstrap_splitting, bachoc2016, tian2018selective) and in particular in the linear model conditioned on LASSO selection (lassoinference, tibshirani2016, liu2018powerful, panigrahi2021integrative). Closest to our work is liu2018powerful, which can be thought of as adjusting the standard $t$-test (or, more accurately, the $z$-test, since they assume $\sigma^{2}$ known) for a coefficient to make it valid conditional on that coefficient being selected by the LASSO. The conditional $\ell$-test can be thought of as making an analogous adjustment to the $\ell$-test rather than the $t$-test, resulting in similar gains under sparsity over liu2018powerful’s method as the unconditional $\ell$-test achieves over the standard $t$-test; see Sections 5, 6.4, and Appendix H.4 for further comparison and discussion.
@@ 1.5 Software
An R (r) package implementing our methods is available at https://github.com/SSouhardya/ell_test_R and scripts for replicating all analyzes in this paper are available at https://github.com/SSouhardya/l-test.
@@ 2.1 The $\ell$-distribution
As the first step towards characterizing the $\ell$-distribution $F^{\lambda}_{\ell}(\cdot\mid\bm{S}^{(j)})$, we restate luo2022 that exactly characterizes the distribution of $\bm{y}\mid\bm{S}^{(j)}$ under $H_{j}$. Let $\bm{P}_{-j}=\bm{X}_{-j}(\bm{X}_{-j}^{T}\bm{X}_{-j})^{-1}\bm{X}_{-j}^{T}$ denote the projection matrix onto the column space of $\bm{X}_{-j}$.
Lemma 1 (luo2022).
@@
where $\mathbb{S}^{n-d}$ denotes the unit sphere of dimension ndn-d.
For completeness, a proof of the lemma is provided in Section D.1 of the Appendix. Next our main theoretical result, Theorem 1, establishes a mapping between $\hat{\beta}_{j}^{\lambda}$ and just the first element of $\bm{u}$, $u_{1}$ (there is nothing special about index 1 here except that we defined $\bm{V}$ to have only its first column non-orthogonal to $\bm{X}_{j}$), which will give an immediate characterization of $F_{\ell}^{\lambda}$ via the known distribution of $u_{1}$ from Equation (2.2).
Theorem 1 (Characterization of the $\ell$-distribution).
Consider the unique decomposition (2.1) from Lemma 1 and for any $\lambda>0$, $b\in\mathbb{R}$ and $\epsilon\in\{-1,1\}$, define the functions
@@
Theorem 1 exactly characterizes the $\ell$-distribution: for $b\neq 0$, we have
@@
Thus, in particular, if we let $F_{u}$ denote the CDF of $u_{1}$ under $H_{j}$ (because the null distribution of $u_{1}$ does not depend on $\bm{S}^{(j)}$, this is also the CDF conditional on $\bm{S}^{(j)}$), then $F_{\ell}^{\lambda}(b)=\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}\leq b\mid\bm{S}^{(j)})=F_{u}(\Lambda_{j}(b,\mathrm{sign}(b)))$ because $\Lambda_{j}(b,\mathrm{sign}(b))$ is a function of $\bm{S}^{(j)}$ for any fixed $b$. Similarly, for $b=0$, it follows from Theorem 1 that $F_{\ell}^{\lambda}(0)=F_{u}\left(\Lambda_{j}(0,1)\right)$. Note that $F_{u}$ is easily evaluated via a one-to-one mapping to a $t$-distribution, namely, $\frac{\sqrt{n-d}u_{1}}{\sqrt{1-u_{1}^{2}}}\sim t_{n-d}$ (see Appendix D.5 for a proof), which, along with the relations above, can be used to explicitly calculate quantiles of the $\ell$-distribution. The proof of Theorem 1 in Appendix D.3 hinges on two main ideas—first, we use block-wise coordinate descent to characterize the event $\{\hat{\beta}_{j}^{\lambda}=b\}$ in terms of $u_{1}$, and second, we characterize $f_{\bm{S}^{(j)}}$ by obtaining an exact expression for $\frac{\partial}{\partial u_{1}}f_{\bm{S}^{(j)}}$, which turns out to be non-negative throughout, thereby showing $f_{\bm{S}^{(j)}}$ is non-decreasing. We will see next that Theorem 1 also provides critical insights into the power of the $\ell$-test.